# Liu and Zhang

### v2

@@ 2.6 Optimality of Index Policy
Theorem 2.13
(Anderson and Nash 1987) Assume $x$ is feasible for (LP) and $y^{*}$ is feasible for (LD), then
@@
For $A:X\rightarrow Y,~{}x^{1}\mapsto Ax^{1}$, define $Ax^{1}(\Omega)=\sum_{\boldsymbol{\omega}\in\Omega}A_{\boldsymbol{\omega}}^{\Omega}x_{\boldsymbol{\omega}}^{1},\Omega\in\tilde{\Omega}$. Let $Y$ be the linear vector space spanned by $b$ and $Ax^{1},~{}x^{1}\in X$. The positive cone is $P_{Y}=\{y\in Y:y(\Omega)\geq 0,\Omega\subset\Omega(\boldsymbol{\omega_{0}}),y(\Omega(\boldsymbol{\omega_{0}}))=0\}$. By Lemma 2.9, $\Omega_{\boldsymbol{\omega_{i}}}=\{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}}):\lambda_{\boldsymbol{\omega_{j}}}\leq\lambda_{\boldsymbol{\omega_{i}}}\}$. Let $\Omega_{\boldsymbol{\omega_{i}}}^{-}:=\{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}}):\lambda_{\boldsymbol{\omega_{j}}}<\lambda_{\boldsymbol{\omega_{i}}}\}$. In infinite dimensional problems, for any $y\in Y$, define $\lambda^{*}\in Y^{*}$ as follows:
@@
Thus $\langle Ax^{W}-b,\lambda^{*}\rangle=0$ holds. The establishment of the second complementary slackness condition requires an important equation that is trivially satisfied in finite-state problems. However, for the case of countable states considered here, the situation is much more complex and we prove it in the theorem below based on the following assumption. {assumption} Sequence $\{\lambda_{\boldsymbol{\omega}}\}_{\boldsymbol{\omega}\in\Omega(\boldsymbol{\omega_{0}})\cup\{0\}}$ has finitely many limit points. In fact, since state update rules are themselves finite and the consideration of stationary policies is often sufficient for long-term optimization problems, the above assumption is naturally satisfied in practice. Based on this assumption, we give the following important equation.
Theorem 2.14
@@
Proof 2.16
Proof. Let $p_{T,ij}^{S}$ be the probability of being in state $j$ after $T$ time steps starting from state $i$ under the $S$-priority policy. Then we have
@@
[…]
To prove Theorem 2.14, we only need to consider the case where the sequence $\{\lambda_{i}\}_{i\in\mathcal{M}}$ has one limit point. For notation convenience, we renumber the states in $\mathcal{M}$ as $\{*,1,2,\cdots\}$. The state indices in the sequence have the following relationship:
@@
Proof 2.18
Proof. We first prove that $\lim\limits_{n\rightarrow+\infty}R_{i}^{\Omega_{2n-1}^{c}}=R_{i}^{\Omega_{*}^{c}}$. We will use induction to show that for any $T$, $\lim\limits_{T\rightarrow+\infty}R_{T,i}^{\Omega_{2n-1}^{c}}=R_{T,i}^{\Omega_{*}^{c}}$. When $T=1$, if $i\notin\Omega_{*}^{c}$, then for any $n$, $R_{1,i}^{\Omega_{2n-1}^{c}}=R_{1,i}^{\Omega_{*}^{c}}=0$; If $i\in\Omega_{*}^{c}$, since $\lambda_{*}$ is the limit point, there $\exists N$ s.t. $i\in\Omega_{2n-1}^{c}$ when $n>N$. In this case, $R_{1,i}^{\Omega_{2n-1}^{c}}=R_{1,i}^{\Omega_{*}^{c}}=R_{i}$. Thus $\lim\limits_{n\rightarrow+\infty}R_{1,i}^{\Omega_{2n-1}^{c}}=R_{1,i}^{\Omega_{*}^{c}}$ holds for every $i\in\mathcal{M}$.
@@
By Theorem 7.11 in Rudin et al. (1976), $\lim\limits_{n\rightarrow+\infty}\lim\limits_{T\rightarrow+\infty}R_{T,i}^{\Omega_{2n-1}^{c}}=\lim\limits_{T\rightarrow+\infty}\lim\limits_{n\rightarrow+\infty}R_{T,i}^{\Omega_{2n-1}^{c}}$. By Lemma 2.15,
@@
[…]
Lemma 2.19
@@
Proof 2.20
Proof. Here we only prove that $\lim\limits_{n\rightarrow+\infty}W_{i}^{\Omega_{2n+1}}=W_{i}^{\Omega_{*}}$, other identities can be proved similarly. By definition,
@@
[…]
Proof 2.21
Proof of Theorem 2.14. We will separately consider three cases where $j$ is odd, $j=*$ and $j$ is even.
@@
(iii) If $j$ is even, we can assume $j=2N$, then by Lemma 2.19, we have:
@@
[…]
Theorem 2.14 is a natural generalization of the key equation in finite-state problems to countable-state problems. Based on Theorem 2.14, we can establish the second relaxed complementary condition:
Lemma 2.22
@@
Proof 2.23
Proof. By Lemma 2.11 and Theorem 2.14, we have
@@
Thus $\langle x^{W},A^{*}\lambda^{*}-R\rangle=0$Halmos
From the proof of Lemma 2.22, we know that $\lambda^{*}$ satisfies $A^{*}\lambda^{*}-\boldsymbol{R}\in P_{X^{*}}$. Therefore, the remaining condition in the weak duality theorem is $-\lambda^{*}\in P_{Y^{*}}$. To prove this statement, we need to show that, $\forall y\in P_{Y},\langle y,~{}\lambda^{*}\rangle\leq 0$. In fact,
@@
Proof 2.25
Proof. It is easy to see that $\Omega_{*}=\lim\limits_{n\rightarrow+\infty}\Omega_{2n-1}$ and $\Omega_{*}^{-}=\lim\limits_{n\rightarrow+\infty}\Omega_{2n}$. Now we will prove that $Ax^{1}(\Omega_{*})\leq\lim\limits_{n\rightarrow+\infty}Ax^{1}(\Omega_{2n+1})$ and $\lim\limits_{n\rightarrow+\infty}Ax^{1}(\Omega_{2n})\leq Ax^{1}(\Omega_{*}^{-})$. First we show that
@@
In fact, by Theorem 2.4, if the initial state is $\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})$, then for any $\Omega\subset\Omega(\boldsymbol{\omega_{0}})$, $b(\{0\}\cup\Omega)=\frac{1}{1-\beta}-T_{j}^{\Omega^{c}}$. As for $b(\Omega)=\inf\limits_{\pi}\{\sum\limits_{\boldsymbol{\omega}}A_{\boldsymbol{\omega}}^{\Omega}x_{\boldsymbol{\omega}}^{1}(\pi)\}$, when we always active the auxiliary arm, then $\sum\limits_{\boldsymbol{\omega}}A_{\boldsymbol{\omega}}^{\Omega}x_{\boldsymbol{\omega}}^{1}=0$. Combining this fact with $b(\Omega)\geq 0$, we have $b(\Omega)=0$. Next, we only need to consider whether state 00 is in $\Omega_{*}$ or not:
@@
Based on the previous discussion and Lemma 2.24, we arrive at:
Lemma 2.26
$-\lambda^{*}\in P_{Y^{*}}$
In summary, when the problem satisfies PCL-indexability and Assumption 2.6 holds, we have the following main theorem:
[…]
If the countable-state RMAB problem considered above is PCL-indexable and Assumption 2.6 holds, then the optimal policy for the two-arm problem is the priority policy given by $\{\lambda_{\boldsymbol{\omega_{i}}}\}_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}$, i.e., for any $\boldsymbol{\omega_{i}},\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})$, if $\lambda_{\boldsymbol{\omega_{i}}}>\lambda_{\boldsymbol{\omega_{j}}}$, then the policy gives higher priority to state $\boldsymbol{\omega_{i}}$ than to state $\boldsymbol{\omega_{j}}$. Furthermore, the problem is Whittle indexable, and $\lambda_{\boldsymbol{\omega_{i}}}$ is exactly the Whittle index of the belief state $\boldsymbol{\omega_{i}}$.
Proof 2.28
[…]
@@ 3.1 $T$-Step Approximate State Space
Definition 3.1 ($T$-step approximate state space)
@@
Extensive numerical simulations give the following two very helpful observations:
Observation 1
@@
Under Observations 1 and 2, we know that if the number of time slots $T$ is sufficiently large, then the new belief states produced after $T$-step updates can be approximated by some belief states in $\tilde{\Omega}(T|\boldsymbol{\omega})$ which is a finite set.
@@ 3.2 AG Algorithm for Approximate State Space
Now we are ready to incorporate the approximate state space $\tilde{\Omega}(T,\varepsilon|\boldsymbol{\omega}_{0})$ into the AG algorithms introduced in Niño-Mora (2002) based on Theorem 2.27 and Observations 1 and 2. Specifically, for the state update, if the resulting belief state $\boldsymbol{\omega}$ does not fall within $\tilde{\Omega}(T,\varepsilon|\boldsymbol{\omega}_{0})$, then we search for the belief state $\boldsymbol{\omega^{\prime}}$ in $\tilde{\Omega}(T,\varepsilon|\boldsymbol{\omega_{0}})$ that has the closest Euclidean distance to $\boldsymbol{\omega}$, and use the Whittle index $\lambda_{\boldsymbol{\omega^{\prime}}}$ of $\boldsymbol{\omega^{\prime}}$ as the Whittle index of $\boldsymbol{\omega}$ or adopt a linear interpolation since $\varepsilon$ is small. The detailed algorithm for computing the approximate Whittle index for our RMAB with general observation models is given below.
@@
Remark. In the adaptive-greedy part of this algorithm, we need to input the parameter $\boldsymbol{A}$ of every arm. For the approximate state space, we can solve for all the parameters $A_{\boldsymbol{\omega}}^{\Omega}$ through the dynamic programming equations of $T_{\boldsymbol{\omega}}^{\Omega}$. The output FAIL value is used to judge whether the problem is PCL-indexable w.r.t. $\tilde{\Omega}$. If FAIL=0, the problem is PCL-indexable, otherwise not.
@@ 3.3 Numerical Results
In this subsection, we demonstrate the strong performance of our algorithm by numerical simulations. For the calculation of approxiamte state space, we set the maximum number of iterations to 6, and the error limit $\varepsilon$ to $31e-$ 3. For each example, we set the number of arms to 6, and the number of arms activated each time to 1, and other parameters to meet the PCL-indexability. We mainly consider the unit-time-return performance of PCL/Whittle index policy and myopic policy when the RMAB system runs for 200 time slots, using 10000 Monte-Carlo simulation experiments. Figures 2-6 show that our PCL/Whittle index policy achieves a superior performance (specific parameter settings are shown in Table 1-3).
Figure 1: Example 1
Figure 2: Example 2
Figure 3: Example 3
Figure 4: Example 4
Figure 5: Example 5
Figure 6: Example 6
Table 1: Transition Matrices Setting 1
Arm System 1 2 3
1 $\textbf{P}^{(1)}=\begin{pmatrix}0.524&0.221&0.255\\ 0.271&0.277&0.452\\ 0.337&0.281&0.382\end{pmatrix}$ $\textbf{P}^{(1)}=\begin{pmatrix}0.427&0.212&0.361\\ 0.305&0.320&0.375\\ 0.089&0.383&0.528\end{pmatrix}$ $\textbf{P}^{(1)}=\begin{pmatrix}0.305&0.333&0.362\\ 0.206&0.331&0.463\\ 0.236&0.179&0.585\end{pmatrix}$
2 $\textbf{P}^{(2)}=\begin{pmatrix}0.671&0.022&0.307\\ 0.396&0.215&0.389\\ 0.095&0.339&0.566\end{pmatrix}$ $\textbf{P}^{(2)}=\begin{pmatrix}0.504&0.172&0.324\\ 0.355&0.315&0.330\\ 0.129&0.303&0.568\end{pmatrix}$ $\textbf{P}^{(2)}=\begin{pmatrix}0.477&0.187&0.336\\ 0.391&0.266&0.343\\ 0.009&0.404&0.587\end{pmatrix}$
3 $\textbf{P}^{(3)}=\begin{pmatrix}0.371&0.237&0.392\\ 0.235&0.491&0.274\\ 0.443&0.154&0.403\end{pmatrix}$ $\textbf{P}^{(3)}=\begin{pmatrix}0.185&0.259&0.556\\ 0.160&0.595&0.245\\ 0.096&0.378&0.526\end{pmatrix}$ $\textbf{P}^{(3)}=\begin{pmatrix}0.255&0.356&0.389\\ 0.232&0.436&0.332\\ 0.175&0.083&0.742\end{pmatrix}$
4 $\textbf{P}^{(4)}=\begin{pmatrix}0.438&0.298&0.264\\ 0.413&0.460&0.127\\ 0.368&0.233&0.399\end{pmatrix}$ $\textbf{P}^{(4)}=\begin{pmatrix}0.575&0.279&0.146\\ 0.384&0.353&0.263\\ 0.068&0.201&0.731\end{pmatrix}$ $\textbf{P}^{(4)}=\begin{pmatrix}0.670&0.081&0.249\\ 0.712&0.102&0.186\\ 0.256&0.531&0.213\end{pmatrix}$
5 $\textbf{P}^{(5)}=\begin{pmatrix}0.610&0.229&0.161\\ 0.183&0.445&0.372\\ 0.565&0.162&0.273\end{pmatrix}$ $\textbf{P}^{(5)}=\begin{pmatrix}0.261&0.308&0.431\\ 0.215&0.709&0.076\\ 0.422&0.278&0.300\end{pmatrix}$ $\textbf{P}^{(5)}=\begin{pmatrix}0.394&0.085&0.521\\ 0.261&0.648&0.091\\ 0.467&0.133&0.400\end{pmatrix}$
6 $\textbf{P}^{(6)}=\begin{pmatrix}0.346&0.500&0.154\\ 0.392&0.414&0.194\\ 0.497&0.231&0.272\end{pmatrix}$ $\textbf{P}^{(6)}=\begin{pmatrix}0.301&0.436&0.263\\ 0.379&0.464&0.157\\ 0.372&0.152&0.476\end{pmatrix}$ $\textbf{P}^{(6)}=\begin{pmatrix}0.538&0.134&0.328\\ 0.700&0.117&0.183\\ 0.380&0.297&0.323\end{pmatrix}$
Table 2: Transition Matrices Setting 2
Arm System 4 5 6
1 $\textbf{P}^{(1)}=\begin{pmatrix}0.410&0.350&0.240\\ 0.412&0.403&0.185\\ 0.473&0.198&0.329\end{pmatrix}$ $\textbf{P}^{(1)}=\begin{pmatrix}0.586&0.261&0.153\\ 0.164&0.468&0.368\\ 0.018&0.395&0.587\end{pmatrix}$ $\textbf{P}^{(1)}=\begin{pmatrix}0.420&0.214&0.366\\ 0.317&0.338&0.345\\ 0.117&0.429&0.454\end{pmatrix}$
2 $\textbf{P}^{(2)}=\begin{pmatrix}0.524&0.221&0.255\\ 0.271&0.277&0.452\\ 0.337&0.281&0.382\end{pmatrix}$ $\textbf{P}^{(2)}=\begin{pmatrix}0.468&0.461&0.071\\ 0.457&0.448&0.095\\ 0.038&0.550&0.412\end{pmatrix}$ $\textbf{P}^{(2)}=\begin{pmatrix}0.285&0.362&0.353\\ 0.289&0.373&0.338\\ 0.037&0.378&0.585\end{pmatrix}$
3 $\textbf{P}^{(3)}=\begin{pmatrix}0.364&0.299&0.337\\ 0.346&0.480&0.174\\ 0.322&0.420&0.258\end{pmatrix}$ $\textbf{P}^{(3)}=\begin{pmatrix}0.263&0.392&0.345\\ 0.317&0.624&0.059\\ 0.247&0.565&0.188\end{pmatrix}$ $\textbf{P}^{(3)}=\begin{pmatrix}0.538&0.172&0.290\\ 0.379&0.468&0.153\\ 0.408&0.439&0.153\end{pmatrix}$
4 $\textbf{P}^{(4)}=\begin{pmatrix}0.383&0.601&0.016\\ 0.158&0.262&0.580\\ 0.537&0.393&0.070\end{pmatrix}$ $\textbf{P}^{(4)}=\begin{pmatrix}0.179&0.284&0.537\\ 0.322&0.674&0.004\\ 0.372&0.186&0.442\end{pmatrix}$ $\textbf{P}^{(4)}=\begin{pmatrix}0.286&0.581&0.133\\ 0.332&0.491&0.177\\ 0.406&0.352&0.242\end{pmatrix}$
5 $\textbf{P}^{(5)}=\begin{pmatrix}0.370&0.154&0.476\\ 0.440&0.178&0.382\\ 0.413&0.231&0.356\end{pmatrix}$ $\textbf{P}^{(5)}=\begin{pmatrix}0.419&0.253&0.328\\ 0.391&0.267&0.342\\ 0.202&0.393&0.405\end{pmatrix}$ $\textbf{P}^{(5)}=\begin{pmatrix}0.320&0.306&0.374\\ 0.341&0.381&0.278\\ 0.206&0.568&0.226\end{pmatrix}$
6 $\textbf{P}^{(6)}=\begin{pmatrix}0.262&0.432&0.306\\ 0.294&0.508&0.198\\ 0.250&0.339&0.411\end{pmatrix}$ $\textbf{P}^{(6)}=\begin{pmatrix}0.387&0.016&0.597\\ 0.270&0.283&0.447\\ 0.136&0.480&0.384\end{pmatrix}$ $\textbf{P}^{(6)}=\begin{pmatrix}0.163&0.102&0.735\\ 0.053&0.529&0.418\\ 0.121&0.435&0.444\end{pmatrix}$
Table 3: Error and Reward Matrix Setting
Error Matrix Reward Matrix
$\textbf{E}=\begin{pmatrix}0.90&0.05&0.05\\ 0.10&0.80&0.10\\ 0.10&0.05&0.85\end{pmatrix}$ $\textbf{R}=\begin{pmatrix}0&0&0\\ 0&1&0\\ 0&0&2\end{pmatrix}$
Note: the same error and reward matrices were set for each arm in all experiments.

### v3

@@ 2.6 Optimality of Index Policy
Theorem 4.
(Anderson and Nash, 1987) Assume $x$ is feasible for (LP) and $y^{*}$ is feasible for (LD), then
@@
For $A:X\rightarrow Y,~{}x^{1}\mapsto Ax^{1}$, define $Ax^{1}(\Omega)=\sum_{\boldsymbol{\omega}\in\Omega}A_{\boldsymbol{\omega}}^{\Omega}x_{\boldsymbol{\omega}}^{1},\Omega\in\tilde{\Omega}$. Let $Y$ be the linear vector space spanned by $b$ and $Ax^{1},~{}x^{1}\in X$. The positive cone is $P_{Y}=\{y\in Y:y(\Omega)\geq 0,\Omega\subset\Omega(\boldsymbol{\omega_{0}}),y(\Omega(\boldsymbol{\omega_{0}}))=0\}$. By Lemma 2, $\Omega_{\boldsymbol{\omega_{i}}}=\{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}}):\lambda_{\boldsymbol{\omega_{j}}}\leq\lambda_{\boldsymbol{\omega_{i}}}\}$. Let $\Omega_{\boldsymbol{\omega_{i}}}^{-}:=\{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}}):\lambda_{\boldsymbol{\omega_{j}}}<\lambda_{\boldsymbol{\omega_{i}}}\}$. In infinite dimensional problems, for any $y\in Y$, define $\lambda^{*}\in Y^{*}$ as follows:
@@
Thus $\langle Ax^{W}-b,\lambda^{*}\rangle=0$ holds. The establishment of the second complementary slackness condition requires an important equation that is trivially satisfied in finite-state problems. However, for the case of countable states considered here, the situation is much more complex and we prove it in the theorem below based on the following assumption.
Assumption 1.
Sequence $\{\lambda_{\boldsymbol{\omega}}\}_{\boldsymbol{\omega}\in\Omega(\boldsymbol{\omega_{0}})\cup\{0\}}$ has finitely many limit points.
Theorem 5.
@@
Proof.
Let $p_{T,ij}^{S}$ be the probability of being in state $j$ after $T$ time steps starting from state $i$ under the $S$-priority policy. Then we have
@@
Since $|\sum\limits_{j\in\mathcal{M}}p_{T,ij}^{S}R_{j}^{S}|\leq\sum\limits_{j\in\mathcal{M}}p_{T,ij}^{S}|R_{j}^{S}|\leq\frac{C}{1-\beta}$, we have $|R_{i}^{S}-R_{T,i}^{S}|\leq\frac{\beta^{T}C}{1-\beta}\rightarrow 0(T\rightarrow+\infty)$. $\Box$
To prove Theorem 5, we only need to consider the case where the sequence $\{\lambda_{i}\}_{i\in\mathcal{M}}$ has one limit point. For notation convenience, we renumber the states in $\mathcal{M}$ as $\{*,1,2,\cdots\}$. The state indices in the sequence have the following relationship:
@@
Proof.
We first prove that $\lim\limits_{n\rightarrow+\infty}R_{i}^{\Omega_{2n-1}^{c}}=R_{i}^{\Omega_{*}^{c}}$. We will use induction to show that for any $T$, $\lim\limits_{T\rightarrow+\infty}R_{T,i}^{\Omega_{2n-1}^{c}}=R_{T,i}^{\Omega_{*}^{c}}$. When $T=1$, if $i\notin\Omega_{*}^{c}$, then for any $n$, $R_{1,i}^{\Omega_{2n-1}^{c}}=R_{1,i}^{\Omega_{*}^{c}}=0$; If $i\in\Omega_{*}^{c}$, since $\lambda_{*}$ is the limit point, there $\exists N$ s.t. $i\in\Omega_{2n-1}^{c}$ when $n>N$. In this case, $R_{1,i}^{\Omega_{2n-1}^{c}}=R_{1,i}^{\Omega_{*}^{c}}=R_{i}$. Thus $\lim\limits_{n\rightarrow+\infty}R_{1,i}^{\Omega_{2n-1}^{c}}=R_{1,i}^{\Omega_{*}^{c}}$ holds for every $i\in\mathcal{M}$.
@@
By Theorem 7.11 in Rudin et al. (1976), $\lim\limits_{n\rightarrow+\infty}\lim\limits_{T\rightarrow+\infty}R_{T,i}^{\Omega_{2n-1}^{c}}=\lim\limits_{T\rightarrow+\infty}\lim\limits_{n\rightarrow+\infty}R_{T,i}^{\Omega_{2n-1}^{c}}$. By Lemma 6,
@@
Using similar arguments, it can be shown that $\lim\limits_{n\rightarrow+\infty}R_{i}^{\Omega_{2n}^{c}}=R_{i}^{\Omega_{*}^{-c}},~{}\forall i\in\mathcal{M}$. $\Box$
Lemma 8.
@@
Proof.
Here we only prove that $\lim\limits_{n\rightarrow+\infty}W_{i}^{\Omega_{2n+1}}=W_{i}^{\Omega_{*}}$, other identities can be proved similarly. By definition,
@@
Thus we have $\lim\limits_{n\rightarrow+\infty}W_{i}^{\Omega_{2n+1}}=W_{i}^{\Omega_{*}}$. $\Box$
Proof of Theorem 5.
We will separately consider three cases where $j$ is odd, $j=*$ and $j$ is even.
@@
(iii) If $j$ is even, we can assume $j=2N$, then by Lemma 8, we have:
@@
It shows that $W_{j}^{\Omega_{j}}=R_{j}+\sum\limits_{i:\lambda_{i}>\lambda_{j}}(W_{j}^{\Omega_{i}}-W_{j}^{\Omega_{i}^{-}})$ and Theorem 5 is proved. $\Box$
Theorem 5 is a natural generalization of the key equation in finite-state problems to countable-state problems. Based on Theorem 5, we can establish the second relaxed complementary condition:
Lemma 9.
@@
Proof.
By Lemma 3 and Theorem 5, we have
@@
Thus $\langle x^{W},A^{*}\lambda^{*}-R\rangle=0$Box
From the proof of Lemma 9, we know that $\lambda^{*}$ satisfies $A^{*}\lambda^{*}-\boldsymbol{R}\in P_{X^{*}}$. Therefore, the remaining condition in the weak duality theorem is $-\lambda^{*}\in P_{Y^{*}}$. To prove this statement, we need to show that, $\forall y\in P_{Y},\langle y,~{}\lambda^{*}\rangle\leq 0$. In fact,
@@
Proof.
It is easy to see that $\Omega_{*}=\lim\limits_{n\rightarrow+\infty}\Omega_{2n-1}$ and $\Omega_{*}^{-}=\lim\limits_{n\rightarrow+\infty}\Omega_{2n}$. Now we will prove that $Ax^{1}(\Omega_{*})\leq\lim\limits_{n\rightarrow+\infty}Ax^{1}(\Omega_{2n+1})$ and $\lim\limits_{n\rightarrow+\infty}Ax^{1}(\Omega_{2n})\leq Ax^{1}(\Omega_{*}^{-})$. First we show that
@@
In fact, by Theorem 1, if the initial state is $\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})$, then for any $\Omega\subset\Omega(\boldsymbol{\omega_{0}})$, $b(\{0\}\cup\Omega)=\frac{1}{1-\beta}-T_{j}^{\Omega^{c}}$. As for $b(\Omega)=\inf\limits_{\pi}\{\sum\limits_{\boldsymbol{\omega}}A_{\boldsymbol{\omega}}^{\Omega}x_{\boldsymbol{\omega}}^{1}(\pi)\}$, when we always active the auxiliary arm, then $\sum\limits_{\boldsymbol{\omega}}A_{\boldsymbol{\omega}}^{\Omega}x_{\boldsymbol{\omega}}^{1}=0$. Combining this fact with $b(\Omega)\geq 0$, we have $b(\Omega)=0$. Next, we only need to consider whether state 00 is in $\Omega_{*}$ or not:
@@
(ii) If $0\notin\Omega_{*}$, then $\exists N$, when $n>N$, $0\notin\Omega_{2n-1}$. In this case, $b(\Omega_{*})=b(\Omega_{2n-1})=0$. In summary, we obtain $b(\Omega_{*})\leq\lim\limits_{n\rightarrow+\infty}b(\Omega_{2n+1})$. Similar arguments show that $\lim\limits_{n\rightarrow+\infty}b(\Omega_{2n})\leq b(\Omega_{*}^{-})$. $\Box$
Based on the previous discussion and Lemma 10, we arrive at:
Lemma 11.
$-\lambda^{*}\in P_{Y^{*}}$
In summary, when the problem satisfies PCL-indexability and Assumption 1 holds, we have the following main theorem:
@@
If the countable-state RMAB problem considered above is PCL-indexable and Assumption 1 holds, then the optimal policy for the two-arm problem is the priority policy given by $\{\lambda_{\boldsymbol{\omega_{i}}}\}_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}$, i.e., for any $\boldsymbol{\omega_{i}},\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})$, if $\lambda_{\boldsymbol{\omega_{i}}}>\lambda_{\boldsymbol{\omega_{j}}}$, then the policy gives higher priority to state $\boldsymbol{\omega_{i}}$ than to state $\boldsymbol{\omega_{j}}$. Furthermore, the problem is Whittle indexable, and $\lambda_{\boldsymbol{\omega_{i}}}$ is exactly the Whittle index of the belief state $\boldsymbol{\omega_{i}}$.
Proof.
From the weak duality theorem and above discussion, we know that the optimal policy is the priority policy given by $\{\lambda_{\boldsymbol{\omega_{i}}}\}$. It follows from $A_{0}^{\{0\}\cup\Omega}=1$ and $W_{0}^{\{0\}\cup\Omega}=\lambda$ that the priority index for state 0 is $\lambda$. Therefore, for state $\boldsymbol{\omega_{i}}$, when the system’s subsidy $\lambda=\lambda_{\boldsymbol{\omega_{i}}}$, state $\boldsymbol{\omega}$ is precisely at the critical point for taking active or passive action under the optimal policy. Therefore, this two-arm problem is Whittle indexable, and the Whittle index for state $\boldsymbol{\omega_{i}}$ is $\lambda_{\boldsymbol{\omega_{i}}}$. $\Box$
@@ 3.1 $T$-Step Approximate State Space
Definition 7 ($T$-step approximate state space).
@@
Fact 1.
@@
Facts 1 and 2 can be easily proved by the compactness of the belief space $\Omega$ (a compact simplex). By these facts, for a sufficiently large number $T$ of time slots, the belief states can be well approximated by a finite set of belief states within $\tilde{\Omega}(T|\boldsymbol{\omega})$.
@@ 3.2 AG Algorithm for Approximate State Space
Now we are ready to incorporate the approximate state space $\tilde{\Omega}(T,\varepsilon|\boldsymbol{\omega}_{0})$ into the AG algorithms introduced in Niño-Mora (2002) based on Theorem 12 and Facts 1 and 2. Specifically, for the state update, if the resulting belief state $\boldsymbol{\omega}$ does not fall within $\tilde{\Omega}(T,\varepsilon|\boldsymbol{\omega}_{0})$, then we look for the belief state $\boldsymbol{\omega^{\prime}}$ in $\tilde{\Omega}(T,\varepsilon|\boldsymbol{\omega_{0}})$ that has the closest Euclidean distance to $\boldsymbol{\omega}$, and use the Whittle index $\lambda_{\boldsymbol{\omega^{\prime}}}$ of $\boldsymbol{\omega^{\prime}}$ as the Whittle index of $\boldsymbol{\omega}$ or adopt a linear interpolation since $\varepsilon$ is small. The detailed algorithm for computing the approximate Whittle index for our RMAB with general observation models is given below.
@@
Remark. In the adaptive-greedy part of this algorithm, we need to input the parameter $\boldsymbol{A}$ of every arm. For the approximate state space, we can solve for all parameters $A_{\boldsymbol{\omega}}^{\Omega}$ using the dynamic programming equations of $T_{\boldsymbol{\omega}}^{\Omega}$. The output FAIL value is used to judge whether the problem is PCL-indexable w.r.t. $\tilde{\Omega}$. If FAIL=0, the problem is PCL-indexable, otherwise not.
@@ 3.3 Numerical Results
In this subsection, we present the robust performance of our algorithm through extensive numerical simulations. To approximate the state space, we set the maximum number of iterations to 6 and the error limit $\varepsilon$ to $31e-$ 3. We expanded our simulation framework to evaluate systems of varying scales, ranging from 3x3 to 6x6 configurations (for the transition matrix of arm states). For each arm state size, we conducted simulations with 30, 40, and 50 arms, while maintaining the activation of a single arm per time slot and system parameters adhered to the PCL-indexability. Our primary focus is on comparing the unit-time-return performance of the PCL/Whittle index policy against a myopic policy over 200 time slots, utilizing 10,000 Monte Carlo simulation experiments for each configuration. The performance gap, quantified as the percentage of improvement from the Whittle index policy over the myopic policy, is summarized in Table 1. As shown in all the figures below, the PCL/Whittle index policy consistently achieves significantly superior performance compared to the myopic policy.
Table 1: Performance Improvement of Whittle Index Policy over Myopic Policy
Arm State Size Number of Arms Myopic Policy’s Average Reward Whittle Index Policy’s Average Reward Performance Gain
3x3 30 2.3618 2.4591 4.12%
3x3 40 2.3740 2.5889 9.05%
3x3 50 2.3587 2.4784 5.07%
4x4 30 2.5201 2.7764 10.17%
4x4 40 2.6595 2.7900 4.91%
4x4 50 2.4355 2.7620 13.41%
5x5 30 2.9509 3.1787 7.72%
5x5 40 2.4096 2.5853 7.29%
5x5 50 2.3188 2.5419 9.62%
6x6 30 2.2340 2.4113 7.93%
6x6 40 2.2554 2.4409 8.22%
6x6 50 2.2977 2.4729 7.63%
(a) 30 arms
(b) 40 arms
(c) 50 arms
Figure 1: Performance comparison for 3x3 system with varying number of arms
(a) 30 arms
(b) 40 arms
(c) 50 arms
Figure 2: Performance comparison for 4x4 system with varying number of arms
(a) 30 arms
(b) 40 arms
(c) 50 arms
Figure 3: Performance comparison for 5x5 system with varying number of arms
(a) 30 arms
(b) 40 arms
(c) 50 arms
Figure 4: Performance comparison for 6x6 system with varying number of arms